given tan(x) = 3/4 0 is less than or equal to x is less then or equal to pi/2 find sin (x)/2
Answer by lwsshak3(11628)

given tan(x) = 3/4 0 is less than or equal to x is less then or equal to pi/2 find sin (x)/2
***
working with a (3-4-5) reference right triangle in quadrant I.
sinx=3/5
sinx/2=3/10